the legs 6 ft and 8 ft what the hypotense
Answer by mananth(16946)

By pythagoras theorem
side1^2+side2^2= hypotenuse ^2
6^2+8^2= hypotenuse^2
36+64=hypotenuse ^2
100 = hyptenuse ^2
square root
hypotenuse = 10 ft